OpenStudy (anonymous):

Select the best possible first step to solving the following system by first eliminating the y variable. 4x + y = 3 5x – 2y = 7 Multiply the first equation by 2. Multiply the second equation by 2. Multiply the first equation by –2. None of the above.

OpenStudy (amistre64):

the "best" step according to who?

OpenStudy (amistre64):

the most useful step might be the first option ....

OpenStudy (anonymous):

wolfram alpha would be my first choice

OpenStudy (anonymous):

Multiply the first equation by 2.

OpenStudy (anonymous):

they want you to say the first one

jhonyy9 (jhonyy9):

the first is right

OpenStudy (anonymous):

By multiplying the first equation by 2, you get 8x + 2y = 6 Then if you add this to the second equation, you get 13x = 13. Therefore, x=1 PLug that into the first equation and you get 4(1)+y=3 5+y=3 y=-2 (1,-2)